solve undetermined coefficient y'' - y' = sinx

1
Expert's answer
2021-05-31T01:59:39-0400

Given equation is -"{y^{''}}-y'=sinx" ......A)


"\dfrac{d^{2}y}{dx^{2}}-\dfrac{dy}{dx}= sinx"


D = "\dfrac{d}{dx}"


writing the equation in normal form -


"(D^{2}-D)y=sinx"


writing in auxiliary form we get ,


auxiliary equation is -


"m^{2}-m =0"


m=0,1


CF can be written as "y=C_1e^{0x}+C_2e^{x}"


"y=C_1+C_2e^{x}"


now , we have to this by undetermined coefficient , for "sinx" method let the particular solution is "y=Acosx+Bsinx"


"y^{'}=-Asinx+Bcosx"


"y{''}=-Acosx-Bsinx"


now putting the value of "y^{'},y{''}" in equation A) , we get -


"=-Acosx-Bsinx +Asinx-Bcosx=sinx"


"=cosx(-A-B)" "+sinx(-B+A)" ="sinx"


comparing the coefficient of "sinx\ and \ cosx \ o"n both side of equation -


"=-A-B=0"

"=A-B=1"

on solving we get ,

"=" "B=\dfrac{-1}{2}"


"=A=\dfrac{1}{2}"


So our particular solution will be -


"y=\dfrac{1}{2}cos x+\dfrac{1}{2}sinx"


now complete solution is given by -


"y=CF+PI"


"y=" "C_1+C_2e^{x}+\dfrac{1}{2}(cosx+sinx)"